A plane flies with a ground speed of 160 mph if there is no wind. It travels 350 miles with a headwind in the same time it takes to go 450 miles with a tailwind. Find the speed of the wind
Answer by mananth(12270)

plane speed = 160 km/h
wind speed = x km/h
wind speed
speed against wind 160 - x
speed with wind 160 + x

Distance against wind 350 km
Distance with wind 450 km

t=d/r time against = time with wind
350/(160-x)=450 /(160+x )
350(160 +x)=450 (160-x)
56000-+ 350 x = 72000 -450 x
450x+350x = 16000
800 x = 16000
/ 800
x = 20 mph wind speed
m.ananth@hotmail.ca